# Keystroke Dynamics: Concepts, Techniques, and Applications

### v2

@@ 4.3 Deep Learning-Based Algorithms
Deep learning algorithms have become the new state of the art in the field of keystroke dynamics. Deep learning is a machine learning technique which learns features automatically from data. The data in this case are sequential keystroke data. Deep learning, when modeled properly, is robust to natural variations in sequential or temporal data. In keystroke dynamics research, usually a conventional machine learning algorithm works as a binary classifier whereas a deep learning algorithm works as a global model that generates embeddings for all subjects.
Deep learning is data hungry and its performance is partially directly proportional to the volume of the training data used, and as such, its application in keystroke dynamics has been limited due to the lack of large publicly available datasets. However, with the availability of large keystroke datasets such as the Aalto [35, 90] and Clarkson II [60] dataset, this option is now being explored [1, 2]. Popular deep neural network types are the convolutional neural network (CNN) and recurrent neural network (RNN).
@@ 4.3.1 Convolutional Neural Network
Convolutional neural networks (CNN’s) [51] are deep learning algorithms that take in an input image and learn to identify key features in the image mich like human vision. The basic CNN architecture is comprised of several layers such as the input layer, convolution layer where high-level features are extracted from the input image, pooling layer where the spatial size of the convolved feature is reduced, and the fully connected or dense layer where the result is flattened out and passed through an activation function that outputs the final classification result [40]. Basically, CNN uses convolution, also known as a sliding window (known as a “filter”) that passes over the image, extracting key features feeding them to lower layers in the network. A CNN, when trained with enough data, can successfully capture the spatial and temporal information in the data. Although CNN was originally created for image classification where pixels form a two-dimensional grid, it has found its usefulness in keystroke dynamics [25, 74, 129] and several other fields.
@@
To leverage the CNN for keystroke analysis, keystrokes data are grouped into sequences, with each sequence containing a fixed length ($L$) of sequential keystroke data. Each sequence is then converted into a two-dimensional grid of specified dimension $L\times B$ (similar to an image), and all sequences are passed as input into the CNN model to be trained, where $L$ is the sequence length (also known as timesteps) and $B$ is the number of features. The overall input size is represented by - [samples ($N$), timesteps ($L$), features ($B$)], where samples ($N$) is the total number of sequences created from the dataset. There are two main types of CNN that can be used - the two-dimensional CNN (2D-CNN) and the one-dimensional CNN (1D-CNN). The 2D-CNN is the standard CNN where the kernel (filter) slides along 2 dimensions on the input data and extracting the spatial features [74]. The 1D-CNN on the other hand, has a kernel that covers all the features and slides along one dimension [76]. Both the 1D-CNN and 2D-CNN have shown to be effective on time-series sequential data such as the keystroke data, however, the 1D-CNN is preferable and has produced better results in literature. This is because 1D-CNNs are designed to handle sequences of one-dimensional vectors, while 2D-CNNs are designed to handle images or data represented in two dimensions. In keystroke dynamics, the data is usually represented as sequences of time-based features including but not limited to hold time and inter-key time. 1D-CNNs can capture the temporal patterns in the sequences of these features, whereas 2D-CNNs would not be well-suited for this type of data. Additionally, due to the sequental characteristic of keystroke data extensive padding would be required to produce a 2D representation, which would be computationally expensive and may negatively affect model performance. 1D-CNNs are computationally more efficient and better suited for keystroke dynamics data, as they can effectively handle the sequential patterns in the data without the need for padding. Although CNN is less computationally intensive as contracted with recurrent neural network (RNN), it can not process all keystroke data in a sequence simultaneously and it is unable to capture the sequence order information like RNN would.
@@ 4.3.2 Recurrent Neural Network
The recurrent neural network (RNN) [78] is a deep learning algorithm best used for sequential or time series data such as keystroke data. A RNN keeps track of information from previous inputs through its memory gate, and decides which information in the past should be remembered or forgotten based on the current input and output. As a result, unlike other deep neural networks, the output of RNN depends on previous information in the sequence, especially for a unidirectional RNN. However, a bidirectional RNN can produce a output that depends on both the past and future information in the sequence.
4.3.2.1 Long Short Term Memory
is a popular RNN architecture, introduced by Hochreiter and Schmidhuber [57]. It is used in modeling long-distance relations in input data. The main benefit of LSTM is inclusion of a cell state, which simply acts like a memory chain by storing information from past states. There are three gates in LSTM - an input gate, an output gate, and a forget gate. The three gates act as logic control within the model used to predict the model’s output. The forget gate is computed using equation $f_{t}=\sigma(W_{f}\cdot[h_{t-1},x_{t}]+b_{f})$; where $\sigma$ is the sigmoid function, $h_{t-1}$ is the output from previous step, $x_{t}$ is the input from current step, and $W_{f}$ and $b_{f}$ are the weight matrix and bias for the forget gate. At the input gate, the computations are carried out with these equations $i_{t}=\sigma(W_{i}\cdot[h_{t-1},x_{t}]+b_{i})$; $\hat{C_{t}}=tanh(W_{c}\cdot[h_{t-1},x_{t}]+b_{c})$. Results are then used to update the cell state using equation $C_{t}=f_{t}\cdot C_{t-1}+i_{t}\cdot\hat{C_{t}}$; where $C_{t-1}$ represents the state of the previous cell. The final output for the current step is calculated using these equations $o_{t}=\sigma(W_{o}\cdot[h_{t-1},x_{t}]+b_{o})$; $h_{t}=o_{t}\cdot tanh(C_{t})$.
@@ 4.3.2.1 Long Short Term Memory
SNN is a class of neural network architectures that contain two or more mirrored sub-networks. A major difference between SNN and other traditional LSTM models is that, a SNN is designed to find the degree of simularity of the inputs to teh sub-networks. Hence, it is able to predict new classes of data from users not seen during training without the need to retrain the model. The architecture achieved state-of-the-art performance in keystroke dynamics authentication with EER of 2.2% and 9.2% on Aalto desktop [35] and mobile [90] datasets, respectively.
4.3.2.2 Gated Recurrent Unit
is also a variant of RNN with similar properties to a LSTM [29]. Here, instead of a “cell state” and three gates used in LSTM, GRU uses hidden states and has two gates - a reset gate and an update gate. The reset and update gates control which information to retain. The GRU is without the memory unit which makes it expose the full hidden content without any control. With fewer gates, GRU trains faster, are less complex and performs better than LSTM, especially on smaller dataset [68, 73]. In relation to keystroke dynamics, [68] contrasted the performance of GRU and LSTM on the same dataset using two scenarios, first was using only the dwell time, and the second was using all the data. They reported that the GRU (15% EER) outperformed LSTM (21.9% EER) when only the dwell time is used, but the LSTM (13.6% EER) outperformed GRU (22.4% EER) when all the date was used.
@@ 4.3.3 Transformer
In [110], Stragapede et al. proposed a novel Transformer architecture named TypeFormer for mobile keystroke dynamics for user authentication purpose. Their proposed Transformer network was based on a two-branch architecture: Temporal and Channel Modules enclosing two Long Short-Term Memory (LSTM) recurrent layers, Gaussian Range Encoding (GRE), a multi-head Self-Attention mechanism, and a Block-Recurrent Transformer layer. TypeFormer could map slices of keystroke sequences into a featute embedding space. Same subject representations of sequences were closer than those of different subjects. The triplet loss function was used to train TypeFormer and the similarity of the feature embeddings was measured with Euclidean distance. The pre-processed Transformer input sequences were fed to the branches. A GRE was used for modelling the input sequences in both branches; the purpose was to preserve the information position. The Temporal Module consisted of three ordered sets of layers. Each of the sets of layers was composed respectively by N, R, and M layers. The N and M layers were made of two sub-layers: a multi-head Self-Attention mechanism, and a multi-scale keystroke LSTM RNN layer. R recurrent layers were included between the N and M layers. The Channel Module consisted of H layers similar to the N and M layers of the Temporal Module. There was a convolutional layer at the end of each of the Modules. The output feature embedding vector P was generated by concatenating the similarity of the output features. The output feature embedding vectors were compared through the Euclidean distance for the authentication task. Stragapede used TypeFormer on the Aalto mobile keystroke database [90] and achieved an EER of 3.25%.

### v3

@@ 4 Keystroke Authentication Algorithms
4.1 Statistical Algorithms
Statistical algorithms are easy to implement, model, and explain. They are deterministic and also do not require heavy computing power and/or memory. However, they require manual extraction of features which could be difficult. Below, we describe common statistical algorithms used for keystroke dynamics.
4.1.1 Gaussian Mixture Model
Gaussian Mixture Model (GMM) is a statistical method based on the weighted sum of probability density functions of multiple Gaussian distributions [122, 34]. GMM generates a vector of mean values corresponding to each component and a matrix of covariance which includes components’ variances and the co-variances between each other. The parameter set $\lambda$ is used for expressing GMM. The parameter $\lambda$ comprises of component weights $w\textsubscript{$i$}$, mean vector $\overrightarrow{\mu\textsubscript{$i$}}$ and covariance matrix $\Sigma\textsubscript{$i$}$ as shown in equation $\lambda=\{w_{i},\overrightarrow{\mu_{i}},\Sigma_{i}\},\text{$i=1,...,M$}$. The parameters are estimated using the iterative expectation–maximization (EM) algorithm [33]. In every iteration, the parameter ($\lambda$) is updated if the iteration yields a higher likelihood and fits the distribution of the training dataset.
Given a set of reference samples $\lambda$ (also known as profile), the mixture density for the keystroke input sample $\overrightarrow{x}$ is defined as the weighted linear combination of M pure Gaussian distributions as shown in the equation: $p(\overrightarrow{x}|\lambda)=\sum_{i=1}^{M}w_{i}*p_{i}(\overrightarrow{x})$; where $p_{i}(\overrightarrow{x})=\frac{1}{(2\pi)^{D/2}|\Sigma_{i}|^{1/2}}exp\biggl{\{}-\frac{1}{2}(\overrightarrow{x}-\overrightarrow{\mu})^{T}\Sigma_{i}^{-1}(\overrightarrow{x}-\overrightarrow{\mu})\biggl{\}}$.
4.1.2 Euclidean Distance
[…]
4.1.3 Manhattan and Scaled Manhattan Distance
[…]
(a) Sample below the median
(b) Sample above the median
Figure 5: Graphical representation of how the ITAD metric is computed from the PDF and CDF. In (a), the sample is below the median and the ITAD metric is determined by the left tail area of the PDF. If the left tail area is x, then the x point of the CDF (on the left) indicates the ITAD metric score. In (b), the sample is above the median and the ITAD metric is determined by the right tail area of the PDF. If the right tail area is y, then the (1-y) point of the CDF (on the right) indicates the ITAD metric score.
4.1.4 Instance-Based Tail Area Density
[…]
The illustration in Figure 5 depicts the computation of the ITAD metric using a graphical representation. The ITAD metric is calculated as the area of the tail of the PDF. If the sample falls below the median, the ITAD metric is determined by the tail area on the left, while if it is above the median, it’s determined by the tail area on the right. ITAD performs better than any distance metrics (such as the Scaled Manhattan) which are based on mean and standard deviation of graphs and are affected by outliers. ITAD metric scores range between 0 and 0.5 and can be interpreted as a measure of similarity. The larger the score is, the closer the test sample is to the profile.
Similarity $\text{Similarity Score}=\frac{1}{N}\sum_{i=1}^{N}S_{\text{$i$}}$ (6)
4.1.5 Mahalanobis Distance
Mahalanobis distance is an example of a distribution-based algorithm. It is an effective multivariate distance metric that measures the distance between a point and a distribution [96, 132, 131, 66]. It takes into consideration the correlation between features which is more suitable when features are not independent. It is best for finding multivariate outliers and finds its applications in multivariate anomaly detection, classification, and one-class classification. The Mahalanobis equation is: $D=(X-\mu_{t})\cdot C^{-1}\cdot(X-\mu_{t})$; where $X$ is the feature vector of a test sample, $\mu_{t}$ is the mean feature calculated from the profile template, and $C^{-1}$ is the inverse covariance matrix of the user profile template.
4.1.6 Gunetti and Picardi’s Metric
Gunetti and Picardi’s free-text algorithm [52] combines typing speed (A-measure) and the degree of disorder (R-measure) to measure similarity. The ‘A’ measure represents the distance between typing samples $S_{1}$ and $S_{2}$ in terms of $n$-graphs, where $n$ is the number of consecutive keystrokes ($n=1$ is monograph, $n$=2 is digraph, $n$=3 is trigraph, etc.), is given in equation -
[…]
The ‘R’ measure on the other hand quantifies the degree of disorder between two sequences. Given an array $V$ of $K$ elements, a simple measure of the degree of disorder of $V$ with respect to $V^{\prime}$ (its ordered counterpart) is computed as the sum of the distances between the position of each element in $V$ and the position of the same element in $V^{\prime}$.
4.2 Machine Learning-Based Algorithms
Machine learning methods have been increasingly employed for keystroke dynamics. Here, we will detail the most commonly used machine learning-based methods that are used for keystroke dynamics.
4.2.1 Supervised Learning
Supervised Machine Learning results in a model that predicts the outcome/label of a given data after the model has been trained with labeled data. The model at training has access to the label of the input (training) data which is used for computing training errors through the loss function. Based on training errors, the model backpropagates and updates the weights. When model training is complete, it is used to predict unseen input data in the absence of labels. Model accuracy is represented by the number of accurate predictions over the total number of predictions.
Supervised Machine Learning can be further divided into classification and regression. Classification uses an algorithm to accurately predict a class label on test data e.g., classifying user behavior as genuine or imposter. Binary classification refers to a problem involving two classes and multi-class classification refers to a problem with greater than three classes. On the other hand, regression makes numerical predictions such as predicting the weather or average salary. In keystroke dynamics, the focus is either to verify if the user is who he/she claimed to be (authentication) or to identify the user to who the query sample belongs to (identification). Both scenarios belong to the classification type, hence, keystroke dynamics uses the classification techniques.
4.2.1.1 Support Vector Machine
A support vector machine (SVM) is a type of supervised learning model leveraged for classification problems. The SVM algorithm computes hyperplane(s) that uniquely separate classes of points of data. The hyperplane, also known as the decision boundary, is an N-dimensional space plane, where N is the number of features in the dataset. Of several hyperplanes possible, the goal is to compute a boundary with the maximum margin that separates the classes (legitimate users from impostors) [46, 69, 128, 107]. This algorithm derives its name from support vectors, which are data points that are closer to the hyperplane and decide the orientation of the plane.
In keystroke dynamics, it is assumed that the keystroke samples in the instance of enrollment and verification have an equal distribution, which is not usually the case, and as a result, lower prediction accuracy is obtained from most traditional machine learning classifiers. In this regard, SVM has support for adaptive learning [26], where an existing SVM classifier (source) is adapted to another SVM classifier (target) based on some projection techniques. Knowing that there exist variations in users’ typing patterns at every sample, especially when the samples have a wide time margin (weeks, months, or years apart), the adaptive SVM modifies the hyperplane of a pre-trained classifier using the new data points with minimum retraining and therefore performs better than the regular SVM trained on the same amount of labeled data. The SVM classifier is a commonly used classifier in keystroke dynamics, especially on a small feature set dataset [24, 44].
4.2.1.2 Naive Bayes
The Naive Bayes classifier is a probabilistic machine learning model used for classification and based on the Bayes theorem with “naive” independent assumptions between features [16, 56, 31]. The Bayes’ theorem finds the probability of an event $a$ occurring given that the probability of another event $B$ has already occurred, as shown in equation $P(a|B)=\frac{P(B|a)\times P(a)}{P(B)}$; where $P(a|B)$ is the posterior probability, $P(a)$ is the prior probability of the label class, $P(B)$ is the prior probability of the predictor and $P(B|a)$ is the likelihood. The Naive Bayes classifier offers several advantages, such as being less computationally demanding compared to other traditional machine learning models, and performing well on small training data when the assumption of feature independence holds. Additionally, it can be used for solving multi-class prediction problems, which are identification problems that involve predicting one of several possible classes (or outcomes) for a given input. In keystroke dynamics, where dependencies between features may exist, the Naive Bayes classifier may not be as effective and is therefore less commonly used.
4.2.1.3 Tree-Based Models
Tree-based models are a type of supervised machine learning algorithm used for classification. They use a set of conditional statements to recursively split training data into subsets. The end result model is a roadmap of logical decisions that describes the dataset. Tree-based models are easy to implement and interpret, less computationally intensive, and are popularly used in keystroke dynamics because of their non-parametric attributes, making it preferable even for non-Gaussian distributed data [104]. Tree-based models are tolerant to outliers and intra-class variation problems, work with both numerical and categorical variables, and require fewer data preprocessing steps such as variable transformations (scaling, normalization, etc). However, they are prone to overfitting and high variance. Popular tree-based methods used in keystroke dynamics are random forest [17, 11], gradient boost and extreme gradient boost (XGBoost) [107, 27]
4.2.1.4 K-Nearest Neighbors
K-Nearest Neighbors (KNN) is a supervised machine learning classifier (although it can also be unsupervised) that classifies new data points based on their proximity with data points of already known classes. KNN works on the assumption that data points of similar classes are closer to each other and can be found near each other [58, 109, 16]. During testing the distances between a new point of data and prior points (neighbors) are calculated. A class label is assigned to a new data point based on a majority vote between its K nearest neighbors, where K is the number of neighbors that will be checked to determine the class to assign for the new query point. Generally, the optimal value of K is found experimentally, as a low value of K results in low bias but high variance, while a high value of K gives high bias and low variance. Distance metrics (such as the Euclidean distance, Manhattan distance, Hamming distance, etc.) are used to calculate the distance between the query point and other data points. These distance metrics form the decision boundaries. KNN is known as a lazy learner as it does the nearest neighbor calculation at every run time. Because of this, it is used only on low-dimensional data and has a high computational cost when used on high-dimensional data.
4.2.1.5 Multi-Layer Perceptron
Multi-layer Perceptron (MLP) is a type of supervised learning algorithm consisting of input, output, and several hidden layers (those in between the input and output layer). The input data to be processed is passed to the input layer, the output layer does the classification, and the hidden layers (non-linear layers) perform the computational work. Therefore, data flows in the forward direction from the input layer to the output layer. MLP is a neural network with multiple layers and each layer is a combination of neurons such that the output of some neurons is passed as input to other neurons in the next layer. The number of neurons and hidden layers used is dependent on the type of classification problem being solved. MLP is capable of learning non-linear models, but it requires several hyperparameter tuning such as the number of hidden neurons, hidden layers, and iterations. Hence, it is computationally intensive and requires both genuine and imposter data to train. The neural network is capable of producing much better performance than statistical algorithms and any other machine learning algorithms when working with big data that has both genuine and impostor samples [5, 53, 113, 92].
4.2.2 Unsupervised Learning
Unsupervised learning is a type of ML used for identifying patterns in datasets that contain unlabeled data points whereas with supervised learning the data is labeled and the algorithm is trained on this data to make predictions. In keystroke dynamics, unsupervised learning algorithms are often used for outlier detection. Some of the unsupervised learning algorithms used in keystroke dynamics include One-class SVM (OCC), K-NN, and K-Means [62, 125]. These algorithms are useful in situations where obtaining labeled data is difficult, costly, or impossible, and the ground truth (label) is unknown.
@@ 4.3 Deep Learning-Based Algorithms
Deep learning algorithms have become the new state-of-the-art in the field of keystroke dynamics. Deep learning is a machine learning technique that learns features automatically from data. The data in this case are sequential keystroke data. Deep learning, when modeled properly, is robust to natural variations in sequential or temporal data. In keystroke dynamics research, usually a conventional machine learning algorithm works as a binary classifier whereas a deep learning algorithm works as a global model that generates embeddings for all subjects.
Deep learning is data-hungry and its performance is partially directly proportional to the volume of the training data used, and as such, its application in keystroke dynamics has been limited due to the lack of large publicly available datasets. However, with the availability of large keystroke datasets such as the Aalto [35, 90] and Clarkson II [60] dataset, this option is now being explored [1, 2]. Popular deep neural network types are the convolutional neural network (CNN) and recurrent neural network (RNN).
@@ 4.3.1 Convolutional Neural Network
Convolutional neural networks (CNNs) [51] are deep learning algorithms that take in an input image and learn to identify key features in the image much like human vision. The basic CNN architecture is comprised of several layers such as the input layer, the convolution layer where high-level features are extracted from the input image, the pooling layer where the spatial size of the convolved feature is reduced, and the fully connected or dense layer where the result is flattened out and passed through an activation function that outputs the final classification result [40]. Basically, CNN uses convolution, also known as a sliding window (“filter”) that passes over the image, extracting key features and feeding them to lower layers in the network. A CNN, when trained with enough data, can successfully capture the spatial and temporal information in the data. Although CNNs were originally created for image classification where pixels form a two-dimensional grid, it has found its usefulness in keystroke dynamics [25, 74, 129] and several other fields.
@@
To leverage the CNN for keystroke analysis, keystroke data are grouped into sequences, with each sequence containing a fixed length ($L$) of sequential keystroke data. Each sequence is then converted into a two-dimensional grid of specified dimension $L\times B$ (similar to an image), and all sequences are passed as input into the CNN model to be trained, where $L$ is the sequence length (also known as timesteps) and $B$ is the number of features. The overall input size is represented by: samples ($N$), timesteps ($L$), and features ($B$) where samples ($N$) is the total number of sequences created from the dataset. There are two main types of CNN that can be used: the two-dimensional CNN (2D-CNN) and the one-dimensional CNN (1D-CNN). The 2D-CNN is the standard CNN where the kernel (filter) slides along 2 dimensions on the input data and extracts the spatial features [74]. The 1D-CNN on the other hand, has a kernel that covers all the features and slides along one dimension [76]. Both the 1D-CNN and 2D-CNN have shown to be effective on time-series sequential data such as the keystroke data, however, the 1D-CNN is preferable and has produced better results in the literature. This is because 1D-CNNs are designed to handle sequences of one-dimensional vectors, while 2D-CNNs are designed to handle images or data represented in two dimensions. In keystroke dynamics, the data is usually represented as sequences of time-based features including but not limited to hold time and inter-key time. 1D-CNNs can capture the temporal patterns in the sequences of these features, whereas 2D-CNNs would not be well-suited for this type of data. Additionally, due to the sequential characteristic of keystroke data extensive padding would be required to produce a 2D representation, which would be computationally expensive and may negatively affect model performance. 1D-CNNs are computationally more efficient and better suited for keystroke dynamics data, as they can effectively handle the sequential patterns in the data without the need for padding. Although CNN is less computationally intensive than a recurrent neural network (RNN), it can not process all keystroke data in a sequence simultaneously and it is unable to capture the sequence order information like RNN would.
@@ 4.3.2 Recurrent Neural Network
The recurrent neural network (RNN) [78] is a deep learning algorithm best used for sequential or time series data such as keystroke data. An RNN keeps track of information from previous inputs through its memory gate and decides which information in the past should be remembered or forgotten based on the current input and output. As a result, unlike other deep neural networks, the output of RNN depends on previous information in the sequence, especially for a unidirectional RNN. However, a bidirectional RNN can produce an output that depends on both the past and future information in the sequence.
Long Short Term Memory is a popular RNN architecture, introduced by Hochreiter and Schmidhuber [57]. It is used to model long-distance relations in input data. The main benefit of LSTM is the inclusion of a cell state, which simply acts like a memory chain by storing information from past states. There are three gates in LSTM: an input gate, an output gate, and a forget gate. The three gates act as logic control within the model used to predict the model’s output. The forget gate is computed using equation $f_{t}=\sigma(W_{f}\cdot[h_{t-1},x_{t}]+b_{f})$; where $\sigma$ is the sigmoid function, $h_{t-1}$ is the output from previous step, $x_{t}$ is the input from current step, and $W_{f}$ and $b_{f}$ are the weight matrix and bias for the forget gate. At the input gate, the computations are carried out with these equations $i_{t}=\sigma(W_{i}\cdot[h_{t-1},x_{t}]+b_{i})$; $\hat{C_{t}}=tanh(W_{c}\cdot[h_{t-1},x_{t}]+b_{c})$. Results are then used to update the cell state using equation $C_{t}=f_{t}\cdot C_{t-1}+i_{t}\cdot\hat{C_{t}}$; where $C_{t-1}$ represents the state of the previous cell. The final output for the current step is calculated using these equations $o_{t}=\sigma(W_{o}\cdot[h_{t-1},x_{t}]+b_{o})$; $h_{t}=o_{t}\cdot tanh(C_{t})$.
@@
SNN is a class of neural network architectures that contain two or more mirrored sub-networks. A major difference between SNN and other traditional LSTM models is that an SNN is designed to find the degree of similarity of the inputs to the sub-networks. Hence, it is able to predict new classes of data from users not seen during training without the need to retrain the model. The architecture achieved state-of-the-art performance in keystroke dynamics authentication with EER of 2.2% and 9.2% on Aalto desktop [35] and mobile [90] datasets, respectively.
Gated Recurrent Unit is also a variant of RNN with similar properties to an LSTM [29]. Instead of a “cell state” and three gates used in LSTM, GRU uses hidden states and has two gates: a reset gate and an update gate. The reset and update gates control which information to retain. The GRU is without the memory unit which makes it expose the full hidden content without any control. With fewer gates, GRU trains faster, is less complex, and performs better than LSTM, especially on smaller datasets [68, 73]. In relation to keystroke dynamics, Kobojek and Saeed [68] contrasted the performance of GRU and LSTM on the same dataset using two scenarios, the first was using only the dwell time, and the second was using all the data. They reported that the GRU (15% EER) outperformed LSTM (21.9% EER) when only the dwell time is used, but the LSTM (13.6% EER) outperformed GRU (22.4% EER) when all the date was used.
@@ 4.3.3 Transformer
Stragapede et al. [110] developed a novel Transformer architecture named TypeFormer for mobile keystroke dynamics for user authentication purpose. Their Transformer network was based on a two-branch architecture: Temporal and Channel Modules enclosing two Long Short-Term Memory (LSTM) recurrent layers, Gaussian Range Encoding (GRE), a multi-head Self-Attention mechanism, and a Block-Recurrent Transformer layer. TypeFormer could map slices of keystroke sequences into a feature embedding space. Same subject representations of sequences were closer than those of different subjects. The triplet loss function was used to train TypeFormer and the similarity of the feature embeddings was measured with Euclidean distance. The pre-processed Transformer input sequences were fed to the branches. A GRE was used for modeling the input sequences in both branches; the purpose was to preserve the information position. The Temporal Module consisted of three ordered sets of layers. Each of the sets of layers was composed respectively by N, R, and M layers. The N and M layers were made of two sub-layers: a multi-head Self-Attention mechanism, and a multi-scale keystroke LSTM RNN layer. R recurrent layers were included between the N and M layers. The Channel Module consisted of H layers similar to the N and M layers of the Temporal Module. There was a convolutional layer at the end of each of the Modules. The output feature embedding vector P was generated by concatenating the similarity of the output features. The output feature embedding vectors were compared through the Euclidean distance for the authentication task. Stragapede used TypeFormer on the Aalto mobile keystroke database [90] and achieved an EER of 3.25%.